# Overestimation of melting temperatures calculated by first-principles molecular dynamics simulations

### v1

@@ Abstract
Although the melting temperature, $T_{m}$, of a solid can be calculated based on first-principles molecular dynamics (FP-MD) simulations, systematic assessments of the accuracy of the resulting values have not yet been reported. FP-MD simulations require significant computational resources and hence an examination of the effect of cell size on convergence is difficult. In addition, calculation of the energy of a liquid is not a trivial problem because of energy dissipation effects. The present work attempts to resolve these problems, and thus allow the accuracy of $T_{m}$ values obtained from FP-MD simulations to be assessed for typical semiconductors, metals, and oxides. With the exception of Si, the $T_{m}$ value was overestimated in all cases. This overestimation can be reduced by increasing the cell size, although the convergence is slow unless the potential is very shallow. For oxides, this overestimation may not be removed by increasing the cell size. The LDA/GGA error of overbinding affects the melting enthalpy and thereby $T_{m}$. In order to fully capture the energy dissipation nature of liquids, adiabatic MD simulations are required, and such simulations have been performed in the present study.
I Introduction
Numerous successes of density-functional theory (DFT) calculations in recent materials research promote the status of DFT calculations to almost an exact science. On this basis, first-principles (FP) molecular dynamics (MD) simulations, which are based on DFT, would be expected to accurately predict the melting temperature ($T_{m}$). Although there are many studies to apply FP-MD simulations to liquids Stich et al. (1991); Kresse and Hafner (1994); Chelikowsky et al. (2001); Massobrio et al. (2015), little studies focus $T_{m}$ as the main topics. The primal interests of these studies are the structural properties of liquids, such as the radial distribution function (RDF), the transport properties, and reaction kinetics. Accordingly, little attention was paid for the accuracy of $T_{m}$ values obtained from FP-MD simulations. Although FP calculations do not involve fitting parameters, control parameters exist in real FP codes and the results are to some extent affected by these parameters. Thus, it is necessary to examine the effects of such parameters on convergence, even though this often requires high computational loads
@@ I Introduction
Many researchers studying melting phenomena continue to use classical MD simulations with model potentials, because of needs for prolong and large-scale simulations. The success of such simulations largely depends on the accuracy of the model potentials, which are usually tested by comparing selected properties of the material between calculation and experiment. As such, structural parameters are most commonly used. Phonon calculations using the force constant model can be looked upon as an example of construction of model potential. Often only few parameters can give a good agreement with an experimental phonon spectrum Bruesch (1982). However, agreement in one case does not guarantee agreement with respect to other properties. As an example, Porter et al. demonstrated that different models that produce almost identical phonon spectra predicted completely different thermal expansion values Porter et al. (1997). This inconsistency could be improved by increasing the number of parameters in a model. Even so, in the case of phenomena for which the harmonic approximation provides an inaccurate prediction, it remain unclear whether a suitable model potential can be constructed based on a reasonable number of parameters. Melting involves increased uncertainty, in that no reliable model superior to the existing Lindemann rule is available Grimvall and Sjödin (1974); Granato et al. (2010).
@@
Reviewing previous studies concerning the melting behavior of silica-based materials illustrates the challenge associated with devising satisfactory models of potential for silica. In fact, there have been many attempts to devise potentials for silica so as to accurately reproduce their physical properties. By constructing interatomic potentials of Si-O pairs in order to match the experimental values, Yamahara et al. were able to correctly reproduce the negative thermal expansion of $\beta$-cristobalite and the $\alpha$-$\beta$ transition of cristobalite of approximately 500 K Yamahara et al. (2001). However, the same potentials produced a $T_{m}$ value for $\beta$-cristobalite about 5000 K, which is much higher than the experimental value of 1996 K. The same MD simulations also largely overestimated for the glass transition temperature, $T_{g}$, giving a value of 3800 K compared with the experimental result of 1480 K. Predicting the glass transition temperature of silica is an especially challenging task. Although many MD simulations have been reported Vollmayr et al. (1996); Yamahara et al. (2001); Kuzuu et al. (2004); Takada et al. (2004); Carré et al. (2008); Geske et al. (2016); Niu et al. (2018), the calculated values of $T_{g}$ are scattered over a wide range of temperatures from 2000 to 4000 K. Takada, et al. report that popular potentials TTM and BKS reproduce the structures of silica well, but provide unsatisfactory $T_{g}$ Takada et al. (2004). Here it is helpful to note that the structural parameters that are matched with experiment when potentials are constructed are the time-averaged structures, such as the bond length. RDF is in this category. The distinction between the solid and liquid states is smeared out in those averaged quantities. The time correlation becomes more important for liquids, as discussed in this paper. We recall the often-claimed leaning that real potentials have highly non-analytic forms, which explains why the DFT approach is so valuable.
@@
DFT should provide, in principle, the correct answer to this problem, because the potentials are constructed from very fundamental principles of physics. However, in practice, use of DFT does not automatically guarantee the correctness of the results. First, even if the potential were exact, the final result will only be correct in the case that the MD simulations are properly set up. The most serious problem in this respect is imposing the periodic-boundary conditions. Melting involves a break-up of the periodicity of crystal, and so melting simulations are performed using the largest supercells that the computational resource permits. In reality, the practical limit is a few times the size of the primitive unit cell. The imposing the periodic-boundary conditions causes several spurious effects. The effects of finite-size supercell are well studied for the defect physics Castleton et al. (2009); Freysoldt et al. (2014). However, different types of consideration are required for the melting phenomenon, which has not been studied in depth to date.
@@
Second, the currently available DFT potentials are still approximations. The exact energy functional in DFT is not known, and will likely not be found in the near future. The overbinding problem of LDA/GGA is a well-known problem with these approximations Jones and Gunnarsson (1989); Parr and Yang (1989). However, the influence of the overbinding on $T_{m}$ is not well studied not (a). This requires calculation of the internal energy of liquid. Although calculation of energy itself is easy in DFT, it is not a trivial task for liquids. Liquids exist only at finite temperatures, and hence the calculation of $E$ without specifying $T$ has no sense. We need to determine both $E$ and $T$ as a set. Thus the problem is equivalent to calculating specific heat $C$, for which the standard theory is lacked for liquids Granato (2002); Trachenko and Brazhkin (2011); Bolmatov et al. (2012); Trachenko and Brazhkin (2016); Proctor (2020); Baggioli and Zaccone (2021). Determination of the relationship between $E$ and $T$ is critical. Energy dissipation (atom relaxation) processes play an important role in determining the $E-T$ relation, and careful consideration of this effect is required for the set up of MD simulation. This is the topic that the present authors have recently established for the glass transition Shirai et al. (2022, 2023). By utilizing the method of calculating $E$ and $T$, we are able to discuss the effect of overbinding on $T_{m}$.
In this paper, the accuracy of $T_{m}$ values obtained by FP-MD simulations is assessed by the method employed in previous studies Shirai et al. (2022, 2023). In Sec. III.1, we report that the $T_{m}$ values obtained by FP-MD simulations are almost always overestimated by examining various materials. The reason for the overestimation is not unique, and depends on material. Further analysis is made for selected materials in the remaining part of Sec. III. These examples are chosen from semiconductors, metals, and oxides. This selection hopefully provides a useful guidance for further study. At present, computationally efficient methods for the determination of $T_{m}$ are available, including thermodynamic integration Frenkel and Smit (1996); Sugino and Car (1995); Cheng et al. (2019). However, the primitive method of adiabatic MD simulations remains important, because the accuracy of calculations is inseparable from an appropriate understanding of the liquid state thermodynamics, as explained in the next section.
II Methods
II.1 Simulation of melting
@@
This work employed both dynamical and thermodynamical methods to determine $T_{m}$ on the basis of MD simulations. The most commonly used method may be the calculation of diffusion constant $D$. $T_{m}$ is determined as the temperature at which a finite value $D$ appears as a solid is heated. The calculation of $D$ is a quite standard method Frenkel and Smit (1996), and thus no explanation is needed here. The other method is the calculation of the internal energy $U$, because $U(T)$ exhibits an abrupt change at $T_{m}$. This, in turn, provides the latent heat associated with the melting enthalpy $H_{m}$. Because of the importance of energetics, we focus more on this $U$ method. Recently, the authors developed a method of calculating the specific heat associated with the glass transition Shirai et al. (2022, 2023). This method is advantageous in that it provides the specific heat and other thermodynamic functions for both solids and liquids with an equal level of accuracy. This method is described below in order to emphasize the full thermodynamic treatment of the liquid state.
II.1.1 Energy vs. temperature
@@
where $E_{\rm gs}(\{{\bf R}_{j}(t)\})$ is the ground state energy of the electrons, including ion-ion potentials, for the instantaneous positions $\{{\bf R}_{j}(t)\}$. The internal energy in thermodynamics is defined at equilibrium and is given by the time average of the microscopic total energy $E_{\rm tot}(t)$ of the system as
@@
where ${\bf D}_{ij}$ is the dynamic matrix between the $i$th and $j$th atoms Born and Huang (1969). The first term of the right-hand side of Eq. (3) is referred to $E_{\rm P,vib}$ herein. The remaining part, $E_{\rm st}(\{\bar{\bf R}_{j}\})$, in the expansion $\overline{E_{\rm gs}(\{{\bf R}_{j}(t)\})}$ is constant with respect to time and is referred to as the structural energy. Thus, the internal energy $U$ consists of the following terms,
@@
Then, the phonon energy $E_{\rm ph}$ is evaluated through Eq. (5). The phonon DOS, $g(\omega)$, is obtained from the power spectrum of atom velocities in an MD run Payne et al. (1992), whereas the temperature in the Bose occupation number is determined by the kinetic energy of atoms via Eq. (6). Finally, $E_{\rm tot}$ is recalculated by adding this phonon energy $E_{\rm ph}$ to $E_{\rm st}$, recovering the Bose-Einstein statistics. In practice, this modification of $E_{\rm tot}$ is important only at very low temperatures.
@@
For liquids, the atoms do not have unique equilibrium positions and hence the expansion (3) does not make sense. Despite this, we can operationally define a “phonon” energy for a liquid in a manner similar to that described above in Eq. (5). This is possible because the frequency spectrum of atom velocities, $f_{v}(\omega)$, can be obtained from MD simulations. In fact, deriving the phonon dispersion in this manner is a practice in the inelastic neutron scattering analysis Egelstaff (1992); Copley and Rowe (1974); Smith et al. (2017). Observation of even a Brillouin-zone-like structure is surprising Suck (2005). The frequency spectrum of atom velocities $f_{v}(\omega)$, however, does not distinguish scattering from vibrational motions. The idea of $f_{v}(\omega)$ applies even to an ideal gas, in that collisions create oscillatory behavior. This phonon-like scenario is often used in modeling liquids, which often gives a good description of the specific heat $C_{p}$ Wallace (1997, 1998); Bolmatov et al. (2012); Trachenko and Brazhkin (2016); Baggioli and Zaccone (2021). We note, however, that the specific heats of many metal liquids are, from beginning, close to the classical limit of the equipartition law of energy, $C_{v}=3k_{\rm B}$ Wallace (2002), which is insensitive to the mode of atom motions. Throughout this paper, the specific heat per atom is understood by using the unit $k_{\rm B}$. Upon closer observation, it will be seen that the phonon-like model described above is only an approximation, as discussed in Sec. III.2. A good example is water, for which $C_{p}\cong 3k_{\rm B}$ even though H-related vibrations cannot be thermally activated at the $T$ range of relevance. Liquid Se has a large $C_{p}$ more than 4 $k_{\rm B}$ Shu et al. (1980). In many metallic liquids, $C_{v}$ near $T_{m}$ falls in a range 3.0 to 3.4 $k_{\rm B}$, which already exceeds the classical limit (Ref. Wallace (2002), p. 242). Some authors ascribe this excess specific heat to the anharmonic effect of phonons Wallace (1998); Bolmatov et al. (2012); Trachenko and Brazhkin (2016); Proctor (2020); Baggioli and Zaccone (2021). But, this excess specific heat increases as $T$ approaches $T_{m}$ from the high temperature side. The anharmonic effect of phonons is expected to be reduced upon lowering $T$ Leibfried and Ludwig (1961). Moreover, recent experiments on supercooled liquids show that the increasing behavior of the excess specific heat as $T$ decreases is continuously extended in the range $T<T_{m}$, namely, $dC_{p}/dT\approx$ const with a negative constant value: Perepezko and Paik (1984) for metallic liquids, Rhim and Ishikawa (2000); Li et al. (2004) for melton Ge. Therefore, it is inappropriate to ascribe the increasing behavior of the excess specific heat with decreasing $T$ to the anharmonic effect of phonons.
@@
In spite of facile success of phonon theory for liquids, the theory eventually fails to accurately describe liquid states not (b). Around the transition temperature of phase transition, the contribution of $E_{\rm st}$ plays an important role in the $T$ dependence on $U$, as shown by previous MD simulations Shirai et al. (2022, 2023). Recent studies on highly viscous glass-forming materials, show that the configuration contribution—this corresponds to $E_{\rm st}$ in this paper—dominates the excess entropy in the supercooled liquid state Smith et al. (2017); Alvarez-Donado and Antonelli (2020); Han et al. (2020). The contribution from $E_{\rm st}$ must therefore be taken into account. In the present method, the total energy $E_{\rm tot}$ includes both the components $E_{\rm st}$ and $E_{\rm ph}$. In the sprit of DFT, the total energy is the most reliable quantity.
@@
The difficulty in calculating the thermodynamic functions of liquids lies essentially in that there is no eigenstate for liquids. In standard applications of statistical mechanics, the presence of eigenstates $\{i\}$ is presumed, when the partition function, $Z=\sum_{i}e^{-\varepsilon_{i}/k_{\rm B}T}$, is evaluated. In the phonon model, phonon-like excitations is envisaged for liquids. But this phonon is soon destroyed after creation. The energy conservation requires that the destruction of a phonon is compensated for by creation of new phonons Leibfried and Ludwig (1961). The liquid state can therefore be considered as a steady state achieved by the dynamic balance between incessant creation and destruction of phonon-like excitations. In an adiabatic simulation the total energy of the system is conserved and the temperature is adjusted to achieve this dynamic balance. Such a system with constant $NVE$ is usually referred to as the microcanonical ensemble, but here the name adiabatic MD simulation is used. The importance of the energy dissipation in determining $T$ is demonstrated by the frequency dependence of the specific heat in the glass transition Birge and Nagel (1985); Nielsen and Dyre (1996); Hentschel et al. (2008). This cannot be explained by the phonon theory. By construction, this energy dissipation process is fully reflected in MD simulations because of the equations of atom motions, and hence the correct $U-T$ relationship can be automatically obtained by MD simulations. However, if the viscose terms coupled to the thermostat are introduced into the MD simulation, the the energy dissipation process is modified by this artificial coupling between the system and the thermostat, destroying the intrinsic relationship between $U$ and $T$. Adiabatic ($NVE$) simulations are therefore necessary to obtain the correct $U-T$ relationship .
II.1.2 Construction of MD simulations

### v2

@@ Abstract
Calculating the melting temperature, $T_{m}$, by molecular dynamics (MD) simulations is a controversial issue. Even first-principles molecular dynamics (FP-MD) simulations often give large overestimations. The accuracy of the $T_{m}$ value obtained by FP-MD simulations has been assessed by examining various solids and calculation parameters. Because the equilibria of liquids are achieved by the balance between many atom relaxation processes, adiabatic simulation—no use of thermostat—is necessary. The present simulations resolved the presence of the width $W_{m}$ for the transition, which vanishes only at the thermodynamic limit. This brings uncertainty to the obtained $T_{m}$ value by $W_{m}$. Within this uncertainty, it has been demonstrated a good agreement in $T_{m}$ is obtained for Si and Na by a reasonable size of cells, namely, approximately 100 atoms. This implies that the contribution of surface in determining $T_{m}$ does not constitute a universal reason for use of the solid-liquid coexistence method for bulk solids, unless too large surface energy or other factors are involved. On the other hand, typically for oxide materials, the FP-MD simulations give large overestimations for $T_{m}$, which are insensitive to the calculation conditions. It is highly likely that the overestimation in $T_{m}$ is brought about by the LDA/GGA error in the electron correlation functional.
1 Introduction
Numerous successes of density-functional theory (DFT) calculations in recent materials research promote the status of DFT calculations to almost an exact science. On this basis, first-principles (FP) molecular dynamics (MD) simulations, which are based on DFT, would be expected to accurately predict the melting temperature ($T_{m}$). In real studies, however, the situation is not simple. Often FP-MD simulations exhibit hysteresis, which makes it difficult to determine uniquely the melting temperature. More seriously, they often give large discrepancies in $T_{m}$, mostly overestimation [1, 2, 3]. The discrepancies depend on the material to be studied. In order to fix the discrepancy, new methods were proposed. Then, the result has the method dependence. This means that the study of a particular material requires the suitable method. This situation is clearly undesirable for material research. To this extent, accurate calculation of $T_{m}$ is so difficult even for DFT.
@@ 1 Introduction
Many researchers studying melting phenomena continue to use classical MD simulations with model potentials, because of needs for prolong and large-scale simulations. The success of such simulations largely depends on the accuracy of the model potentials, which are usually tested by comparing selected properties of the material between calculation and experiment. As such, structural parameters are most commonly used. Phonon calculations using the force constant model can be looked upon as an example of construction of model potential. Often only few parameters can give a good agreement with an experimental phonon spectrum [4]. However, agreement in one case does not guarantee agreement with respect to other properties. As an example, Porter et al. demonstrated that different models that produce almost identical phonon spectra predicted completely different thermal expansion values [5]. This inconsistency could be improved by increasing the number of parameters in a model. Even so, in the case of phenomena for which the harmonic approximation provides an inaccurate prediction, it remain unclear whether a suitable model potential can be constructed based on a reasonable number of parameters. Melting involves increased uncertainty, in that no reliable model superior to the existing Lindemann rule is available [6, 7].
@@
Reviewing previous studies concerning the melting behavior of silica-based materials illustrates the challenge associated with devising satisfactory models of potential. There have been many attempts to devise potentials for silica so as to accurately reproduce their physical properties. By constructing interatomic potentials of Si-O pairs in order to match the experimental values, Yamahara et al. were able to correctly reproduce the negative thermal expansion of $\beta$-cristobalite and the $\alpha$-$\beta$ transition of cristobalite of approximately 500 K [8]. However, the same potentials produced a $T_{m}$ value for $\beta$-cristobalite about 5000 K, which is much higher than the experimental value of 1996 K. The same MD simulations also largely overestimated for the glass transition temperature, $T_{g}$, giving a value of 3800 K compared with the experimental result of 1480 K. Predicting the glass transition temperature of silica is an especially challenging task. Although many MD simulations have been reported [9, 8, 10, 11, 12, 13, 14], the calculated values of $T_{g}$ are scattered over a wide range of temperatures from 2000 to 4000 K. Takada, et al. report that popular potentials TTM and BKS reproduce the structures of silica well, but provide unsatisfactory $T_{g}$ [11]. Belonoshko expresses his surprise as “it is puzzling why the interaction model of Matsui is successful in producing $PVT$ properties and solid-solid transformations but fail to give proper results for the pressure dependence of the melting temperature” [2]. Here it is helpful to note that the structural parameters that are matched with experiment when potentials are constructed are the time-averaged structures, such as the bond length. RDF is in this category. The distinction between the solid and liquid states is smeared out in those averaged quantities. The time correlation becomes more important for liquids, as discussed in this paper. We recall the often-claimed leaning that real potentials have highly non-analytic forms, which explains why the DFT approach is so valuable.
@@
DFT should provide, in principle, the correct answer to this problem, because the potentials are constructed from very fundamental principles of physics. However, in practice, use of DFT does not automatically guarantee the correctness of the results. First, even if the potential were exact, the final result will only be correct in the case that the MD simulations are properly set up. The most serious problem in this respect is imposing the periodic-boundary conditions. Melting involves a break-up of the periodicity of crystal, and thus the imposing the periodic-boundary conditions causes various spurious effects. Such effects are well studied for the defect physics: reviews are given by, for example, [15, 16]. However, different types of consideration are required for the melting phenomenon, which has not been well studied to date. A study by Hong and van de Walle is an example of such a few exceptions [17]. Based on their analysis on the size dependence, they proposed a new method to remove the hysteresis and the overheating problem in small system size.
@@
Second, the currently available DFT potentials are still approximations. The exact energy functional in DFT is not known, and will likely not be found in the near future. The overbinding problem of LDA/GGA is a well-known problem with these approximations [18, 19]. However, the influence of the overbinding on the melting temperature is not well studied [20]. This requires calculation of the internal energy of liquid. Although calculation of energy itself is easy in DFT, it is not a trivial task for liquids. Liquids exist only at finite temperatures, and hence the calculation of $E$ without specifying $T$ has no sense. We need to determine both $E$ and $T$ as a set. Thus the problem is equivalent to calculating specific heat $C$, for which the standard theory is lacked for liquids [21, 22, 23, 24, 25, 26]. Determination of the relationship between $E$ and $T$ is critical. Energy dissipation (atom relaxation) processes important in determining the $E-T$ relation, and careful consideration of this effect is required for the set up of MD simulation. This is the topic that the present authors have recently established for the glass transition [27, 28]. By utilizing the method of calculating $E$ and $T$, we are able to discuss the effect of overbinding on $T_{m}$.
As noted at the beginning, the standard method of FP-MD simulation sometimes fails to give reasonable $T_{m}$ values. In order to fix this problem, several methods were proposed, depending on what cause is envisaged as the discrepancy in $T_{m}$. These include thermodynamic integration [29, 30, 31, 32], the solid-liquid (two-phase) coexistence method [1, 2, 33, 3, 34, 17, 35, 36], and the Z method [37, 38]. By observing that melting is initiated from the surface of a solid, the overestimation of $T_{m}$ in MD simulations is often ascribed to the lack of surface in the standard setup of MD simulation. The two-phase coexistence method was introduced to suppress the superheating due to the lack of surface. Certainly, the surface has influence on determination of $T_{m}$ [39]. An evident example is the reduction of $T_{m}$ in small-size systems, such as gold nanoparticles [40]. However, we need to distinguish superheating of solid from supercooling of liquid. These two processes are not in the relation of reversal processes, which are discussed seriously [41, 42, 43, 37, 38]. For supercooling of liquid, the mechanism of supercooling is well described by the standard theory of crystal nucleation. The surface energy $\gamma_{s}$ creates an energy barrier in the free energy $\Delta G^{\ast}$ for the crystal nucleation, which develops as $\Delta G^{\ast}=const\ \gamma_{s}^{3}(T_{m}/H_{m})^{2}/\Delta T_{m}^{2}$, $H_{m}$ is the melting enthalpy: for example, see [44] (p. 192). The crystal nucleation mostly occurs at surface, which is known as inhomogeneous nucleation. In this sense, the two-phase coexistence method is useful to help the creation of surface nucleation in practically realizable times of simulation. In contrast, for superheating of solid, it is not clear what the substance of energy barrier is. There has been, so far, no direct evidence for the “liquid nucleation”. By using the two-phase coexistence method, Belonoshko demonstrated a reduction in the calculated $T_{m}$ by more than 1000 K for MgO [2]. It is suspicious that the surface barrier has such a large effect on $T_{m}$ for bulk solids; the supercooling can be normally achieved with the reduction $\Delta T_{m}$ by less than 10 % of $T_{m}$.
Generally speaking, the bonding environments of the surface atoms are different from that of bulk atoms. Accordingly, it is reasonable to observe the surface effect for small-size systems, such as nanoparticles. The surface effect can be scaled with the inverse of the system size $L$, which presents the ratio of the number of surface atoms ($N_{S}$) to the number of bulk atoms ($N_{V}$). With a suitably chosen $L$, the reduction of the melting temperature, $\Delta T_{m}$, is given by
$\frac{\Delta T_{m}}{T_{m}}\propto\frac{1}{L}\propto\frac{N_{S}}{N_{V}}.$ (1)
See Eq. (14) in [40]. In view of this scaling, the current size of the two-phase coexistence method is of the same order of the magnitude that nanoparticles have. This raises the doubt that the reduction in $T_{m}$ by the coexistence method could be due to the size effect. If this is the case, the result of the coexistence method turns to be underestimation for the bulk. We need to discuss the overestimation in calculations separately from the superheating of experiment. We should not overlook the possibility that the overestimation in $T_{m}$ is due to calculation errors caused by, for example, use of the small cell size or use of inaccurate potentials. In any respect, neither the mechanism of superheating nor the assessment of the coexistence method is the purpose of the present study. Here, we are interested only in the accuracy of $T_{m}$ in the standard setup of FP-MD simulations.
In this paper, the accuracy of $T_{m}$ values obtained by FP-MD simulations is assessed by the method employed in previous studies [27, 28]. In Sec. 2, the characteristics of the method is explained. It is emphasized that the energy-dissipation processes are essential in determining $T_{m}$. In Sec. 3.1, we report the $T_{m}$ values obtained by FP-MD simulations for various materials. The result shows that the overestimation in $T_{m}$ often occurs. The reason for the overestimation is not unique and depends on material. Further analysis is made for selected materials in the remaining part of Sec. 3. These examples are chosen from semiconductors, metals, and oxides. This selection is quite arbitrary, and there is no reason that the accuracy is in accordance with the material classification. In Sec. 4, the present results are summarized.
2 Methods
2.1 Simulation of melting
@@
This work employed both dynamical and thermodynamical methods to determine $T_{m}$ on the basis of MD simulations. The most commonly used method may be the calculation of diffusion constant $D$. $T_{m}$ is determined as the temperature at which a finite value $D$ appears as a solid is heated. The calculation of $D$ is a quite standard method [31], and thus no explanation is needed here. The other method is the calculation of the internal energy $U$, because $U(T)$ exhibits an abrupt change at $T_{m}$. This, in turn, provides the latent heat associated with the melting enthalpy $H_{m}$. Because of the importance of energetics, we focus more on this $U$ calculation. Recently, the authors developed a method of calculating the specific heat associated with the glass transition [27, 28]. This method is advantageous in that it provides the specific heat and other thermodynamic functions for both solids and liquids with an equal level of accuracy. This method is described below in order to emphasize the full thermodynamic treatment of the liquid state.
2.1.1 Energy versus temperature
@@
where $E_{\rm gs}(\{{\bf R}_{j}(t)\})$ is the ground state energy of the electrons, including ion-ion potentials, for the instantaneous positions $\{{\bf R}_{j}(t)\}$. The internal energy in thermodynamics is defined at equilibrium and is given by the time average of the microscopic total energy $E_{\rm tot}(t)$ of the system,
@@
where ${\bf D}_{ij}$ is the dynamic matrix between the $i$th and $j$th atoms [45]. The first term of the right-hand side of Eq. (4) is referred to $E_{\rm P,vib}$ herein. The remaining part, $E_{\rm st}(\{\bar{\bf R}_{j}\})$, in the expansion $\overline{E_{\rm gs}(\{{\bf R}_{j}(t)\})}$ is constant with respect to time and is referred to as the structural energy. Thus, the internal energy $U$ consists of the following terms,
@@
Then, the phonon energy $E_{\rm ph}$ is evaluated through Eq. (6). The phonon DOS, $g(\omega)$, is obtained from the power spectrum of atom velocities in an MD run [46], whereas the temperature in the Bose occupation number is determined by the kinetic energy of atoms via Eq. (7). Finally, $E_{\rm tot}$ is recalculated by adding this phonon energy $E_{\rm ph}$ to $E_{\rm st}$, recovering the Bose-Einstein statistics. In practice, this modification of $E_{\rm tot}$ is important only at very low temperatures.
@@
For liquids, the atoms do not have unique equilibrium positions and hence the expansion (4) does not make sense. Despite this, we can operationally define a “phonon” energy for a liquid in a manner similar to that leads to Eq. (6). This is because the frequency spectrum of atom velocities, $f_{v}(\omega)$, can be obtained from MD simulations. In fact, deriving the phonon dispersion in this manner is a practice in the inelastic neutron scattering analysis [47, 48, 49]. Observation of even a Brillouin-zone-like structure is surprising [50]. The frequency spectrum of atom velocities $f_{v}(\omega)$, however, does not distinguish scattering from vibrational motions. The idea of $f_{v}(\omega)$ can be applied even for an ideal gas, in that collisions create oscillatory behavior. This phonon-like scenario is often used in modeling liquids, which often gives a good description of the specific heat $C_{p}$ [51, 52, 23, 24, 26]. We note, however, that the specific heats of many metal liquids are, from beginning, close to the classical limit of the equipartition law of energy, $C_{v}=3k_{\rm B}$ [53], which is insensitive to the mode of atom motions. Throughout this paper, the specific heat per atom is understood by using the unit $k_{\rm B}$. Upon closer observation, it will be seen that the phonon-like model described above is only an approximation, as discussed in Sec. 3.2. A good example is water, for which $C_{p}\cong 3k_{\rm B}$ even though H-related vibrations cannot be thermally activated at the $T$ range of relevance. Liquid Se has a large $C_{p}$ more than 4 $k_{\rm B}$ [54]. In many metallic liquids, $C_{v}$ near $T_{m}$ falls in a range 3.0 to 3.4 $k_{\rm B}$, which already exceeds the classical limit (Ref. [53], p. 242). Some authors ascribe this excess specific heat to the anharmonic effect of phonons [52, 23, 24, 25, 26]. But, this excess specific heat increases as $T$ is lowered to $T_{m}$ from the high temperature side. The anharmonic effect of phonons is expected to be reduced upon lowering $T$ [55]. Moreover, recent experiments on supercooled liquids show that the increasing behavior of the excess specific heat as $T$ decreases is continuously extended in the range $T<T_{m}$, namely, $dC_{p}/dT\approx$ const with a negative constant value: [56] for metallic liquids, [57, 58] for melton Ge. Therefore, it is unrealistic to ascribe the increasing behavior of the excess specific heat with decreasing $T$ to the anharmonic effect of phonons.
@@
In spite of facile success of phonon theory for liquids, the theory eventually fails to accurately describe liquid states [59]. Around the transition temperature of phase transition, the contribution of $E_{\rm st}$ plays the predominant role in the $T$ dependence on $U$, as shown by previous MD simulations [27, 28]. Recent studies on highly viscous glass-forming materials, show that the configuration contribution—this corresponds to $E_{\rm st}$ in this paper—dominates the excess entropy in the supercooled liquid state [49, 60, 61]. The contribution from $E_{\rm st}$ must therefore be taken into account. In the present method, the total energy $E_{\rm tot}$ includes both the components $E_{\rm st}$ and $E_{\rm ph}$. In the sprit of DFT, the total energy is the most reliable quantity.
@@
The difficulty in calculating the thermodynamic functions of liquids lies essentially in that there is no eigenstate for liquids. There is no microscopic steady state. In standard applications of statistical mechanics, the presence of eigenstates $\{i\}$ is presumed, when the partition function, $Z=\sum_{i}e^{-\varepsilon_{i}/k_{\rm B}T}$, is evaluated. In the phonon model for liquids, phonon-like excitations are envisaged. But this phonon is soon destroyed after creation. The energy conservation requires that the destruction of a phonon is compensated for by creation of new phonons [55]. The liquid state can therefore be considered as a macroscopic steady state achieved by the dynamic balance between incessant creation and destruction of phonon-like excitations. In an adiabatic simulation in a sense that no heat bath is involved, the total energy of the system is conserved and the temperature is adjusted to achieve this dynamic balance. The importance of the energy dissipation in determining $T$ is demonstrated by the frequency dependence of the specific heat in the glass transition [62, 63, 64]. By construction, this energy dissipation process is fully described by MD simulations because of the rigorous equations of atom motions, and hence the correct $U-T$ relationship can be automatically obtained by MD simulations. However, if the viscose term coupled to the thermostat is introduced into the MD simulation, the energy dissipation processes are modified by this artificial coupling between the system and the thermostat, destroying the intrinsic relationship between $U$ and $T$. The melting temperature is determined by subtle balance among many energy dissipation processes. Therefore adiabatic simulations are essential to obtain the correct $U-T$ relationship.[65] In the glass transition, the effect of atom relaxation is well known to yield hysteresis and complicates the analysis particularly when temperature is scanned in caloric methods. To remove the effect of relaxation, the adiabatic calorimetry is used [66, 67]. To emphasize this feature, we use the term adiabatic MD simulation in place of the widely-used term $NVE$ simulation. How adiabatic methods separate the hysteresis from the time dependence properties are discussed in [68].
2.1.2 Construction of MD simulations